USE THE FOLLOWING DATA TO ANSWER QUESTIONS 35 – 40

Annual Income Statements

2017

2018

Sales

525,800

559,700

COGS

????

149,600

Gross profit

381,000

410,100

Oper. exp

236,100

250,100

Depreciation

15,000

16,160

Operating profit

129,900

143,840

Interest exp.

9,500

10,600

EBT

120,400

133,240

Taxes

34,910

42,880

Net Income

85,490

90,360

Annual Balance Sheets

2017

2018

Cash

440,200

468,600

Accounts rec

278,400

252,300

Inventories

293,500

315,700

Current Assets

1,012,100

1,036,600

Net fixed assets

1,125,000

???

Land

100,200

120,600

Total Assets

2,237,300

2,357,400

Notes payable

250,000

177,500

Accounts payable

110,200

115,800

Accruals

12,900

10,600

Current portion of long-term debt

100,200

195,200

Current Liabilities

473,300

499,100

Long-term debt

1,000,000

900,000

Common stock @ ($0.25 par)

25,000

30,300

Additional paid in capital

500,000

610,200

Retained earnings

239,000

317,800

Total Liabilities & Equity

2,237,300

2,357,400

39. Cash flow from investing in 2018 was $__________.

40. Total dividends paid in 2017 was $________.
Answer #1

39)

Cash flow from investing = Capex = Change in net fixed assets + Depreciation  

Net Fixed Assets in 2018 = Total Assets - ( CA + Land) = 2,357,400 -(1,036,600 + 120,600) = 1,200,200

Cash flow from investing = (1,200,200 - 1,125,000) + 16,160 = $ 91,360

40)

Total dividends paid in 2017 = Net Income in 2017 - Change in retained earnings = 85,490 - (317,800- 239,000) = $ 6,690